OpenStudy (anonymous):

Solve quadratic equation by factoring problem. 14x^3+19x^2+5x=8x

OpenStudy (anonymous):

I got half way: x(14x^2+19x-3)=0

OpenStudy (anonymous):

14x^3 + 19x^2 + 5x = 8x 14x^3 + 19x^2 - 3x = 0 x(14x^2 + 19x - 3) = 0 14x^2 + 19x - 3 = 0 Going to have to solve it with the equation x = (-b ± √b^2 - 4ac)/2a x = (-19 ± √19^2 - 4 x 14 x -3)/2 x 14 x = (-19 ± √361 - 168)/28 x = (-19 ± √193)/28 Not sure if its right, thats what I would do :). Hope this has helped in some way.

OpenStudy (anonymous):

I figured it put, the answer was x=0, -3/2, 1/7. Your process was sort of right, thank you for your help though :-)

OpenStudy (anonymous):

out*

OpenStudy (anonymous):

divide both sides by x. then solve.